A 3.0 kg black ball travels to the right at a speed of 1.7 m / s and collides with a stationary 6.0 kg red ball. After the collision, the black ball bounces and retraces its steps at a speed of 0.50 m / s. What is the speed in m / s of the red ball after the collision?


1
Expert's answer
2021-01-11T11:59:05-0500

Let's write the law of conservation of linear momentum:


"m_1v_{1i}+m_2v_{2i}=-m_1v_{1f}+m_2v_{2f}."

From this equation we can find the speed of the red ball, "v_{2f}":


"v_{2f}=\dfrac{m_1v_{1i}+m_2v_{2i}+m_1v_{1f}}{m_2},""v_{2f}=\dfrac{3.0\ kg\cdot 1.7\ \dfrac{m}{s}+6.0\ kg\cdot 0\ \dfrac{m}{s}+3.0\ kg\cdot 0.50\ \dfrac{m}{s}}{6.0\ kg}=1.1\ \dfrac{m}{s}."


The sign plus means that the red ball moves to the right.

Answer:

"v_{2f}=1.1\ \dfrac{m}{s}."